# DARTS-GT: Differentiable Architecture Search for Graph Transformers with Quantifiable Instance-Specific Interpretability Analysis

### v1

@@ IV Experimental Results
    Quantitative interpretability analysis using our novel framework to assess model transparency. Our interpretability framework evaluation operates at two levels.
@@ IV-B Comparison Against Standard Graph Transformers
Table II presents the performance comparison of DARTS-GT with the standard Graph Transformer (Vanilla-GT) introduced by Dwivedi et al. [19], which forms the foundation of many subsequent works [16]. Unlike recent approaches that enhance graph transformers through parallel MPNN paths [6], novel positional encoding [7], or structural encoding-augmented attention [8], DARTS-GT alters the core attention mechanism through asymmetric query-key value computation and heterogeneous per-layer GNN selection. This is a significant departure from standard GT where Q, K, V share the same computational source without GNNs. This experiment quantifies the improvement of DARTS-GT over the baseline it extends. For rigorous comparison, we standardized most experimental conditions using the GPS configuration [6]. We extend our implementation from their opensource repository, adopting their fixed positional encoding, hidden dimensions, attention heads, dropout rates, and pre/post-processing components for both models. Only the layer count was adjusted to maintain a comparable parameter budget range. This controlled setup (Supp-Tables II,III,IV), prioritizing methodological clarity over peak performance, ensures that observed improvements stem from design changes rather than hyperparameter tuning or auxiliary components.
@@ IV-D Dissecting DARTS-GT’s Architecture Discovery Patterns
Figure 2 visualizes the GNN distributions discovered by DARTS across eight datasets, revealing prominent diversity in architectural choices. The selection patterns range from highly specialized (MolPCBA with 81% GINE dominance) to nearly uniform distributions (ZINC, MolHIV with balanced GNN usage). This demonstrates that DARTS adapts to dataset-specific requirements rather than converging to a universal solution, as expected from a neural architecture search (NAS) algorithm. Interestingly, node-level tasks (PATTERN, CLUSTER) favored GatedGCN. We could also observe the complete absence of GATV2 in Peptides datasets, coupled with GatedGCN’s dominance (87.5% in Peptides-Func). However, these proportions only show that the GNN selection is utilised, but the layer-wise positioning of each GNN, which DARTS optimizes through gradient-based search, is equally critical. A dataset might use 50% GINE, but whether these appear in early or late layers inherently impacts feature extraction and performance (examples of a few searched architectures are given in Supp-Figure 2).
@@ IV-E Ablation Studies: Assymetric vs Symmetric attention
The results obtained consistently favored the DARTS-GT asymmetric design over the symmetric variant across all tested configurations. This validates our architectural choice: decoupling queries from structural encoding ($Q_{m}$ from features, $K_{m}$$V_{m}$ from GNNs) provides superior performance compared to the seemingly natural approach of deriving all attention components from the same GNN source.
@@ IV-F1 Quantitative Interpretability Metrics
Table VII presents the interpretability metrics derived from our causal ablation framework (Section III-C). Through systematic head masking on each test instance, we compute head deviations that quantify causal importance, then aggregate using median values for robustness to outliers. The median Specialization captures the typical concentration of head importance across instances, while median Focus measures consensus among important heads on critical nodes. Supp-Figures 4 and 5 display population-level distributions as scatter plots, with median values indicated, for all tests under consideration. We evaluated on two diverse benchmarks: MolHIV (OGB, molecular graphs, classification) and Pep-Struc (LRGB, peptide graphs, regression), representing different domains, graph sizes, and task types. For transparency, we report GPS [6] with both sparse attention (matching our implementation) and dense attention (authors’ original) on MolHIV, revealing how attention mechanism choices significantly impact interpretability ( Focus: 0.5734 sparse vs 0.111 dense).
@@
These metrics characterize the complexity of the interpretation rather than establishing absolute superiority. Our evaluation on two datasets provides initial empirical evidence on real-world applications but, like any empirical study, cannot guarantee trends generalize across all benchmarks. However, one of our primary contributions, the interpretability framework itself, enables practitioners to quantify the difficulty of interpretation. Instance-level analysis (Supp-Figures 4 and 5) reveals considerable variation across graphs, reminding us that specific instances may present unique interpretability challenges regardless of model-level trends.
@@
Figure 3: Instance analysis of Graph 1 (Peptides-Struc, 209 nodes) demonstrating deviation profiles and top-3 attention heatmaps.
@@ IV-G Instance-Specific Interpretation Analysis
Figure 3 presents a representative instance from the Peptides-Struc datasets, where deviation profiles identify truly important heads regardless of visual prominence and attention heatmaps of these top-k heads reveal critical graph regions. Each attention head is uniquely identified by its layer and head index, following the format L<layer_index>_H<head_index>. For example, L0_H0 refers to the first head in the first layer. While we demonstrate Peptides-struct example here, MolHIV is moved to supplementary (Supp-Figure 3) for space constraints.
@@
To our knowledge, the above is the first instance-specific quantitative method for Graph Transformers to identify which heads and nodes causally drive predictions. We have quantitatively shown further (Supp-Figure 1, Supp-Section I), that entropy heuristics cannot replace our head-deviation based causal ablation method. Although space constraints limit us to two representative examples (additional instances available in our GitHub repository), these were strategically selected: Graph 1 (Peptides-struct) showcases a deliberate paradox where visual prominence do not always correlate with causal importance while Graph 50 (MolHIV) presented in supplementary section III, demonstrates our framework’s standard operation and how these metrics varies based on attention types.
Across instances, we observed no reliable correlation between attention salience and causal head importance. In GPS, the most critical head (L0_H0, $\delta=0.317$) appears sparse/dark while colorful heads contribute minimally. In contrast, in Vanilla-GT, a visually prominent dark pattern (L2_H1) proves causally less important than a uniform/colorful head (L0_H0, $\delta=0.262$). This instance-specific variability, where concentrated attention may be causally critical, irrelevant, or anywhere between, demonstrates that practitioners cannot determine head importance from visualization alone. The quantitative deviation analysis of our framework resolves this ambiguity by directly measuring the casual contribution of each head through ablation, regardless of visual appearance. This enables reliable identification of functionally significant components that attention visualization would miss or misrepresent. Users can generate similar analyses for any instance using our code provided, extending beyond these illustrative examples to their specific use cases.
@@ V Conclusion and Future Directions
Addressing Q1 - Architectural Performance and Adaptability: Our experiments definitively demonstrate that Graph Transformers can achieve superior performance through dynamic, layer-wise GNN selection. DARTS-GT consistently outperformed vanilla Graph Transformers on all benchmarks, with relative improvements ranging from 0.7% to 318% depending on the complexity of the dataset. Against state-of-the-art methods, DARTS-GT achieved the best new results on 4 of 8 datasets (ZINC, MalNet-Tiny, Peptides-Func, and Peptides-Struc), while remaining competitive on others.
@@
Future Directions: The convergence of improved performance and quantifiable interpretability opens exciting possibilities for scientific discovery. An appropriate real-world impactful domain application to test our DARTS-GT with the interpretability framework would be the protein structure-to-function prediction problem, where given a protein with known functions $[F1,F2,F3,F4]$, DARTS-GT could not only predict these functions accurately, but also identify which structural motifs causally contribute to each function. When the model makes a confident prediction on a novel protein, our interpretability framework can distinguish whether the prediction relies on established structural patterns (suggesting genuine functional similarity) or spurious correlations (indicating potential overfitting). For instance, if our framework reveals that predictions for enzymatic activity consistently focus on validated active site regions across diverse proteins, this provides evidence that DARTS-GT has learned biologically meaningful patterns rather than dataset artifacts. This capability could accelerate drug discovery by identifying previously unknown functional sites or revealing why certain mutations disrupt protein function.
@@ VI Acknowledgments
This research was funded by Google DeepMind, with S.C. being supported by Google DeepMind Scholarship for her DPhil studies. We would also like to thank Dr.Petar Veličković for helpful discussions. For the purpose of Open Access, the authors have applied a CC BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission. The authors would like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work. https://doi.org/10.5281/zenodo.22558.

### v2

@@ IV Experimental Results
    Quantitative interpretability analysis using our novel framework to assess model transparency. Our evaluation of the interpretability framework operates on two levels.
@@ IV-B Comparison Against Standard Graph Transformers
Table II presents the performance comparison of DARTS-GT with the standard Graph Transformer (Vanilla-GT) introduced by Dwivedi et al. [19], which forms the foundation of many subsequent works [16]. Unlike recent approaches that enhance graph transformers through parallel MPNN paths [6], novel positional encoding [7], or structural encoding-augmented attention [8], DARTS-GT alters the core attention mechanism through asymmetric query-key value computation and heterogeneous per-layer GNN selection. This is a significant departure from standard GT where Q, K, V share the same computational source without GNNs. This experiment quantifies the improvement of DARTS-GT over the baseline it extends. For rigorous comparison, we standardized most experimental conditions using the GPS configuration [6]. We extend our implementation from their opensource repository, adopting their fixed positional encoding, hidden dimensions, attention heads, dropout rates, and pre/post-processing components for both models. Only the layer count was adjusted to maintain a comparable parameter budget range. This controlled setup (Supplementary-Tables II,III,IV), prioritizing methodological clarity over peak performance, ensures that observed improvements stem from design changes rather than hyperparameter tuning or auxiliary components.
@@ IV-D Dissecting DARTS-GT’s Architecture Discovery Patterns
Figure 2 visualizes the GNN distributions discovered by DARTS across eight datasets, revealing prominent diversity in architectural choices. The selection patterns range from highly specialized (MolPCBA with 81% GINE dominance) to nearly uniform distributions (ZINC, MolHIV with balanced GNN usage). This demonstrates that DARTS adapts to dataset-specific requirements rather than converging to a universal solution, as expected from a neural architecture search (NAS) algorithm. Interestingly, node-level tasks (PATTERN, CLUSTER) favored GatedGCN. We could also observe the complete absence of GATV2 in Peptides datasets, coupled with GatedGCN’s dominance (87.5% in Peptides-Func). However, these proportions only show that the GNN selection is utilised, but the layer-wise positioning of each GNN, which DARTS optimizes through gradient-based search, is equally critical. A dataset might use 50% GINE, but whether these appear in early or late layers inherently impacts feature extraction and performance (examples of a few searched architectures are given in Supplementary Figure 2).
@@ IV-E Ablation Studies: Assymetric vs Symmetric attention
The results obtained consistently favored the DARTS-GT asymmetric design over the symmetric variant in all configurations tested. This validates our architectural choice: decoupling queries from structural encoding ($Q_{m}$ from features, $K_{m}$$V_{m}$ from GNNs) provides superior performance compared to the seemingly natural approach of deriving all attention components from the same GNN source.
@@ IV-F1 Quantitative Interpretability Metrics
Table VII presents the interpretability metrics derived from our causal ablation framework (Section III-C). Through systematic head masking on each test instance, we compute head deviations that quantify causal importance, then aggregate using median values for robustness to outliers. The median Specialization captures the typical concentration of head importance across instances, while median Focus measures consensus among important heads on critical nodes. Supplementary Figures 4 and 5 show population-level distributions as scatter plots, with median values indicated, for all tests under consideration. We evaluated on two diverse benchmarks: MolHIV (OGB, molecular graphs, classification) and Pep-Struc (LRGB, peptide graphs, regression), representing different domains, graph sizes, and task types. For transparency, we report GPS [6] with both sparse attention (matching our implementation) and dense attention (authors’ original) on MolHIV, revealing how attention mechanism choices significantly impact interpretability ( Focus: 0.5734 sparse vs 0.111 dense).
@@
These metrics characterize the complexity of the interpretation rather than establishing absolute superiority. Our evaluation on two datasets provides initial empirical evidence on real-world applications but, like any empirical study, cannot guarantee trends generalize across all benchmarks. However, one of our primary contributions, the interpretability framework itself, enables practitioners to quantify the difficulty of interpretation. Instance-level analysis (Supplementary Figures 4 and 5) reveals considerable variation across graphs, reminding us that specific instances may present unique interpretability challenges regardless of model-level trends.
@@
Figure 3: Instance analysis of Graph 1 (Peptides-Struc, 209 nodes) demonstrating deviation profiles and top-3 attention heatmaps.In the attention heatmaps, columns represent target nodes that receive attention.
@@ IV-G Instance-Specific Interpretation Analysis
Figure 3 presents a representative instance from the Peptides-Struc datasets, where deviation profiles identify truly important heads regardless of visual prominence, and attention heatmaps of these top-k heads reveal critical graph regions. Each attention head is uniquely identified by its layer and head index, following the format L<layer_index>_H<head_index>. For example, L0_H0 refers to the first head in the first layer. While we demonstrate Peptides-struct example here, MolHIV is moved to supplementary (Supplementary Figure 3) for space constraints.
@@
To our knowledge, the above is the first instance-specific quantitative method for Graph Transformers to identify which heads and nodes causally drive predictions. We have quantitatively shown further (Supplementary Figure 1, Supplementary Section I), that entropy heuristics cannot replace our head-deviation based causal ablation method. Although space constraints limit us to two representative examples (additional instances available in our GitHub repository), these were strategically selected: Graph 1 (Peptides-struct) showcases a deliberate paradox where visual prominence do not always correlate with causal importance while Graph 50 (MolHIV) presented in supplementary section III, demonstrates our framework’s standard operation and how these metrics varies based on attention types.
Across instances, we did not observe a reliable correlation between attention salience and causal head importance. In GPS, the most critical head (L0_H0, $\delta=0.317$) appears sparse/dark while colorful heads contribute minimally. In contrast, in Vanilla-GT, a visually prominent dark pattern (L2_H1) proves causally less important than a uniform/colorful head (L0_H0, $\delta=0.262$). This instance-specific variability, where concentrated attention may be causally critical, irrelevant, or anywhere between, demonstrates that practitioners cannot determine head importance from visualization alone. The quantitative deviation analysis of our framework resolves this ambiguity by directly measuring the casual contribution of each head through ablation, regardless of visual appearance. This enables reliable identification of functionally significant components that attention visualization would miss or misrepresent. Users can generate similar analyses for any instance using our code provided, extending beyond these illustrative examples to their specific use cases.
@@ V Conclusion and Future Directions
Addressing Q1 - Architectural Performance and Adaptability: Our experiments demonstrate that Graph Transformers can achieve superior performance through dynamic, layer-wise GNN selection. DARTS-GT consistently outperformed vanilla Graph Transformers on all benchmarks, with relative improvements ranging from 0.7% to 318% depending on the complexity of the dataset. Against state-of-the-art methods, DARTS-GT achieved the best new results on 4 of 8 datasets (ZINC, MalNet-Tiny, Peptides-Func, and Peptides-Struc), while remaining competitive on others.
@@
Future Directions: The convergence of improved performance and quantifiable interpretability opens exciting possibilities for scientific discovery. An appropriate real-world impactful domain application to test our DARTS-GT with the interpretability framework would be the protein structure-to-function prediction problem, where given a protein with known functions $[F1,F2,F3,F4]$, DARTS-GT could not only predict these functions accurately, but also identify which structural motifs causally contribute to each function. When the model makes a confident prediction on a novel protein, our interpretability framework can distinguish whether the prediction relies on established structural patterns (suggesting genuine functional similarity) or spurious correlations (indicating potential overfitting).
@@ VI Acknowledgments
This research was funded by Google DeepMind, with S.C. being supported by Google DeepMind Scholarship for her DPhil studies. The authors also thank Dr.Petar Veličković for helpful discussions. For the purpose of Open Access, the authors have applied a CC BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission. The authors acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work. https://doi.org/10.5281/zenodo.22558.